Rewrite the equation so that y is a function of x.
6y - 6x = 15

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

6y-6x=15
6y=6x+15
y=x+2.5

Answer by stanbon(75887)

Rewrite the equation so that y is a function of x.
6y - 6x = 15
------------------------
Add 6x to both sides to get:
6y = 6x + 15
Divide both sides by 6 to get:
y = x + 5/2
===================
Cheers,
Stan H.